OpenStudy (anonymous):

CHECK MY ANSWER Use the quadratic formula to find the zeros of the function. y = 16x2 + 40x + 25 is the answer {-0.8} ??

OpenStudy (wolf1728):

That is not the answer I get

OpenStudy (wolf1728):

[ -40 ± Sq Root ( 1,600 - 4•16•25 ) ] / 32 which equals [ -40 ± Sq Root ( 0 ) ] / 32 x=-40 / 32 x = -5 / 4